# Whetten Maison Parcollet Dinarelli Estève Avignon UniversitéFrance Thales SIXFrance Samsung AI Center CambridgeUnited Kingdom Université Grenoble AlpesFrance

### v1

@@ 4.4 Additional observations
One other finding is that the correlations between inertia and downstream performance are positive except in the case of out-of-domain ASR at 200k where all the models converged to very similar performance. This indicates that the lower the inertia, the better the clustering, and thus better performance. On the other hand, the correlations with the DB index are negative for ASR indicating the worse the clustering, the better the performance. We believe that this because the DB index penalizes clusters that are close to their nearest neighbors. However, overlapping clusters do not necessarily indicate poor performance; instead, close neighboring clusters suggest efficient use of the latent space.
@@ 5 Discussion
We propose using clustering and rank measures as efficient, unsupervised methods for evaluating the performance of SSL speech models and show that these measures are better indicators of downstream performance than the pre-training loss.
To illustrate the potential compute savings, suppose one pre-trains all the models in this experiment fully (i.e. 400k steps), with our settings this would take around 16,800 GPU hours. If one instead trained all models to 50k and then used the rank to select and train the top 5 models to 400k steps, this would take approximately 6k hours, saving approximately 10,800 GPU hours. We believe that these results can potentially save researchers and engineers much training time particularly when working with in-domain data.

### v2

@@ 4.4 Additional observations
One other finding is that the correlations between inertia and downstream performance are positive except in the case of out-of-domain ASR at 200k steps of pre-training where all the models converged to very similar performance. This indicates that the lower the inertia, the better the clustering, and thus better performance. On the other hand, the correlations with the DB index are negative for ASR indicating the worse the clustering, the better the performance. We believe that this because the DB index penalizes clusters that are close to their nearest neighbors. However, overlapping clusters do not necessarily indicate poor performance; instead, close neighboring clusters suggest efficient use of the latent space.
@@ 5 Discussion
We propose using clustering and rank measures as efficient, unsupervised methods for evaluating the performance of SSL speech models and show that these measures are better indicators of downstream performance than the pre-training loss. To illustrate the potential compute savings, suppose one pre-trains the 30 models from this work to 400k steps. With our settings this would take around 16,800 GPU hours. If one instead pre-trained all models to 50k and then used the rank to select and train only the top 5 models to 400k steps, this would take approximately 6,000 hours, saving approximately 10,800 GPU hours. We believe that future work could involve exploring these methods further and developing more robust measures to accurately evaluate SSL speech models early on in training. We believe that future work could involve exploring these methods further and developing.
6 Acknowledgements
This work received funding from the French ANR E-SSL project N°ANR-22-CE23-0013 and used HPC resources from GENCI–IDRIS projects AD011014732 and A0131013821.